New Leader Actions in Lords32

Ryan Powers

August 26, 2001

1 Basic Concepts Behind New Actions

With the introduction of magic new leader actions will be made available. Many of these will use the leaders Magic Rating. This rating will usually vary from 1 to B just like the military rating. Certain leaders (Wizards are one) will have a minimum higher than 1.

Each action entry follows a certain form. It looks like this:

1.1 New Action Name

Code The letter code for this action.

BAC Base Action Cost, the base number of AP’s that must be spent on this action.

Stat The leader’s stat this action depends on.

Results What this action does.

2 Magical Leader Actions

Here are the new leader actions that relate to the magic system.

2.1 Magical Operation

Code MO BAC 8+ Stat Magic Results This action allows the leader to generate MOC and

MOB points. Each 8 AP spent can be used as a point of MOC or MOB. A leader may only produce 1 MOC point, but may produce as many MOB points as their AP supply allows. The Bonus points may be applied to an operation conducted using the MOC point generated, or may be applied to any other operations performed in the turn.

2.2 Magical Research

Code MR BAC 5+ Stat Magic Results This allows a leader to supervise magical research. A

leader is not necessary, but one with a high magic rating will provide bonuses to the research Here is how magical research is performed. This action allows the leader to spend zero or more MFP on magical research this turn Their are four types of research: General, Question, Class, Specific. For details on research, see the Magical Research document.

2.3 Cast Spell

Code CS BAC as spell Stat Magic Results Allows the leader to cast a spell. Spells are described

in their own section below.

3 Leader Spells

Leader Spells are similar to the Magical Operations but are of a more limited scale. They must be researched just like operations. Below is a small selection of spells which can be considered common knowledge. (They still must be researched, but can be researched by name) Each spell has a AP cost, and a MFP cost to cast. They also have a Magic Level that must be reached in the region they are cast in in order to be possible.

3.1 Glib Tongue

Magic Level 1

AP 2

MFP 1+

Results This spell provides a bonus to the leader’s diplomacy rating from when it is cast till the end of the turn. More than 1 MFP may be spent, with extra MFP providing a bonus to success, and possibly to the level of a successful result.

3.2 Haste

Magic Level 1

AP 2

MFP 4+

Result THis spell provides the caster with 4 additional AP per year after it is cast and until the end of the turn, including the year it is cast in. It only affects a lone leader, joining with troops cancels the effect of this spell. Additional MFP may be paid to increase the chances of success. Beware that this will age the caster 1 additional year per year spent at higher speed. It can be cancelled at any time by the caster.

3.3 Shadow Walk

Magic Level 1

AP 2

MFP 2+

Result This makes the caster harder to detect, giving bonuses on moving in secret and espionage actions performed by the leader. As usual more MFP points increase the chance of success as well as the potential level of success. Can only affect a single leader, not a leader and troops.

3.4 Loyalty

Magic Level 1

AP 2

MFP 1+

Result This is cast upon another leader and improves their loyalty rating for the purposes of revolt checks only. More MFP increases the chances of success as well as the potential success levels.

3.5 Flame Darts

Magic Level 1

AP 0 (see descition)

MFP 1+

Result This is a simple combat spell. It is cast in conjunction with a combat action, and does not take any additional AP beyond the action it is supporting. Each MFP spent provides a more powerful effect on the filed of battle. If not cast in conjunction with a combat, it takes 1AP to cast. IT is primarily meant as a personal combat spell, so do not expect miraculous effects on the field of battle. But yo never know when it will be just that extra edge needed.